A certain sum amounts to Rs.3825 in 4 yrs and to Rs.4050 in 6 yrs.Find the rate% and the sum

Grade:8

1 Answers

Subhash
6 years ago
Let Sum = x and Rate of Interest= y%
therefore, x+(x*y*4/100)=3825 …....(1)
and          x+(x*y*6/100)=4050 …....(2)
by subs. eq 1 from eq. 2, we get--
6xy/100 – 4xy/100 = 225
therefore, 2 xy =22500
xy=11250 ….....(3)
by placing the value of xy into eq. 1
x + 11250*4/100 = 3825
i.e. x+ 450 = 3825
therefore, x = 3825-450 = 3375 ….(4)
by placing value of x (eq. 4) , we get
3375*y=11250
therefore, y=11250/3375 = 10/3
therefore, y=10/3%
Rate of interest = 10/3% or 3.33%